# Ferdinand Schlatt $0$-0002-6032-909X Friedrich-Schiller-Universität Jena Maik Fröbe $0$-0002-1003-981X Friedrich-Schiller-Universität Jena  and  Matthias Hagen $0$-0002-9733-2890 Friedrich-Schiller-Universität Jena

### v1

@@ Abstract.
A wide range of transformer-based language models have been proposed for information retrieval tasks. However, fine-tuning and inference of these models is often complex and requires substantial engineering effort. This paper introduces Lightning IR, a PyTorch Lightning-based framework for fine-tuning and inference of transformer-based language models for information retrieval. Lightning IR provides a modular and extensible architecture that supports all stages of an information retrieval pipeline: from fine-tuning and indexing to searching and re-ranking. It is designed to be straightforward to use, scalable, and reproducible. Lightning IR is available as open-source: https://github.com/webis-de/lightning-ir.
@@ 1. Introduction
Pre-trained transformer-based language models have become a cornerstone in information retrieval (IR) research (Lin et al., 2022). Many different architectures have been proposed, each with their own model implementation and training procedure. This plethora makes fine-tuning different model architectures cumbersome and comparing different model architectures even more so. However, all these models use the same basic building blocks, can be fine-tuned in the same way, and have only minor differences in their inference procedure.
To unify the usage of transformer-based language models in IR we present the Lightning IR framework. Lightning IR builds on and extends PyTorch Lightning (Falcon and team, 2024) to provide several key features that set it apart from existing libraries for neural IR: (1) It is backbone agnostic, i.e., (almost) any HuggingFace (Wolf et al., 2020) transformer-based language model can be used. (2) It supports the entire pipeline of IR models, from fine-tuning and indexing to searching and re-ranking. (3) It is flexible and supports, for example, multi-vector or sparse bi-encoder models and pointwise or listwise cross-encoder models out-of-the-box. (4) It provides an easy-to-use API and CLI for fine-tuning and inference. (5) It is easily configurable, allowing for reproducible experiments and easy model comparison.
In this paper, we compare Lightining IR to existing frameworks, we describe its features and API, and we demonstrate Lightning IR’s capabilities in a series of reproducibility experiments.
@@ 2. Comparison to Similar Frameworks
Several existing frameworks support the development and use of retrieval models, but they differ in the supported stages of the retrieval pipeline and the types of models they can handle (see Table 1 for an overview). Frameworks like RAGatouille, SentenceBERT (Reimers and Gurevych, 2019), or Seismic (Bruch et al., 2024) focus on specific model types. In contrast, frameworks like Experimaestro-IR (Piwowarski, 2020) or PyTerrier (Macdonald et al., 2021) are more generic and support a wide range of models and stages, but each model type requires a separate implementation, and not all model types are available for all pipeline stages.
Lightning IR is a framework that implements models as generic, configurable, and extensible modules. In essence, a model is defined by its backbone encoder model and how the contextualized embeddings of a query and document are combined to compute a relevance score. This design allows Lightning IR to support a wide range of model types. In addition, as all model types use the same API, the same fine-tuning and inference procedures can be used for all model types. This makes it easy to compare different model types and to experiment with new model types.
@@ 3.1. Model
Lightning IR supports two types of models: cross-encoders and bi-encoders. Both model types are built on top of a backbone encoder model from HuggingFace (Wolf et al., 2020). For example, the following code shows how to initialize a new bi-encoder and cross-encoder model using some pre-trained model from HuggingFace, where {HF_MODEL} is a placeholder for the model name.
@@
A configuration class defines how each model type uses the contextualized embeddings generated by the backbone model to compute a relevance score. For example, a ColBERT-style (Khattab and Zaharia, 2020) model does not pool the contextualized embeddings. It instead uses late-interaction to compute relevance scores over the contextualized query and document token embeddings. A ColBERT-style model can be initialized in Lightning IR as follows.
@@
For usability and reproducibility, we combine a Lightning IR model and tokenizer in a PyTorch Lightning module (Falcon and team, 2024). The module is responsible for handling training and inference logic, but it also provides convenience functions to quickly score queries and documents. For example, the following code snippet shows how to compute scores using a pre-trained BERT-based bi-encoder and an ELECTRA-based cross-encoder.
@@ 3.2. Dataset
Lightning IR tightly integrates with ir_datasets (MacAvaney et al., 2021) to provide access to a wide range of common information retrieval datasets, but custom datasets are also supported. Datasets are split into four different classes: Document, Query, Tuple, and Run datasets. A Document dataset iterates over a document collection and is used for indexing. Query datasets iterate over queries and are used for retrieval. Tuple datasets are used for fine-tuning as they contain samples consisting of a query and multiple documents. Run datasets contain ranked documents for a query (optional: relevance judgments) and are used for re-ranking; also fine-tuning is possible by sampling n-tuples from rankings. The following code snippet shows how to load the MS MARCO passage dataset (Nguyen et al., 2016) and the TREC Deep Learning 2019 passage ranking dataset (Craswell et al., 2019).
@@
Multiple datasets can be combined into a single PyTorch Lightning datamodule. Similar to the model’s Lightning module, the datamodule makes fine-tuning and inference easier by handling data sampling and batching. It also cleanly separates fine-tuning and evaluation data. The following code snippet shows how to create a datamodule for fine-tuning a bi-encoder model on MS MARCO triples and evaluating it on the TREC Deep Learning 2019 and 2020 passage ranking datasets.
@@
RunDataset("msmarco-passage/trec-dl-2019"),
@@ 3.3. Trainer
To use the index for retrieval, the path of the index must be passed to a searcher class that matches the indexer class used to create the index. If a dataset has relevance judgments in ir_datasets and evaluation metrics are specified in the module’s configuration, the trainer will automatically evaluate the retrieval effectiveness. The following code snippet shows how to retrieve documents for a query using the indexed documents.
@@ 3.4. CLI
Lightning IR provides a command-line interface (CLI) to simplify the usage of the framework. The CLI is built on top of PyTorch Lightning’s CLI and provides commands for fine-tuning, indexing, searching, and re-ranking. All options are configurable via command-line arguments or a configuration yaml file. The configuration yaml files are especially useful for reproducibility. For example, to fine-tune a bi-encoder model on the MS MARCO triples dataset, the following command can be used.
@@
The CLI also supports indexing, searching, and re-ranking via the index, search, and re_rank commands. The configuration file and command for indexing the MS MARCO passage collection using a fine-tuned bi-encoder model is shown below. The configuration files for searching and re-ranking are similar.
@@ 4. Lightning IR: Supported Models
Lightning IR supports fine-tuning and running inference on a wide range of bi- and cross-encoder models. We have additionally added support for a number of popular models not natively fine-tuned in Lightning IR. This includes all bi-encoders from the sentence transformers library (Reimers and Gurevych, 2019), SPLADE models released by Naver labs (Formal et al., 2021), the official ColBERT checkpoints (Khattab and Zaharia, 2020), and monoT5 and RankT5 models (Nogueira et al., 2020; Zhuang et al., 2022). Additional models can be easily added by providing the corresponding configuration files and adding the model to the Lightning IR model registry.
@@
# class_path: BiEncoderModule
@@
monoELECTRA Large (Schlatt et al., 2024) 0.750 0.791
@@ 5. Reproducibility Experiment
Table 3 compares the effectiveness of our fine-tuned models with the official checkpoints [^5] https://huggingface.co/sentence-transformers/msmarco-bert-base-dot-v5, [^6] https://huggingface.co/naver/splade-v3, [^7] https://huggingface.co/colbert-ir/colbertv2.0 provided by the authors on the TREC 2019 and 2020 Deep Learning track data (Craswell et al., 2019, 2020). Our fine-tuned models achieve competitive effectiveness compared to the official checkpoints. Minor differences between our results and the official checkpoints can be attributed to randomness in the training process and are not statistically significant, except for the SPLADE models (our SPLADE model is more effective on TREC Deep Learning 2019 and less effective on TREC Deep Learning 2020). These results demonstrate that one is capable of reproducing the effectiveness of state-of-the-art models with minimal effort using Lightning IR.
@@ 6. Conclusion
We have introduced Lightning IR, a PyTorch Lightning-based framework that enables straightforward fine-tuning and inference of transformer-based language models for information retrieval. By conducting a reproducibility experiment, we have demonstrated the simplicity and flexibility of the framework’s API and the capabilities of Lightning IR. We were able to fine-tune and evaluate a variety of models that achieve competitive effectiveness compared to state-of-the-art neural ranking models with short code snippets and minimal effort. Future work includes extending Lightning IR to support additional efficient dense, sparse, and multi-vector indexing and retrieval pipelines, such as PLAID (Santhanam et al., 2022) and Seismic (Bruch et al., 2024), to improve the latency and scalability of multi-vector and sparse retrieval models.

### v2

@@ Abstract.
A wide range of transformer-based language models have been proposed for information retrieval tasks. However, including transformer-based models in retrieval pipelines is often complex and requires substantial engineering effort. In this paper, we introduce Lightning IR, an easy-to-use PyTorch Lightning-based framework for applying transformer-based language models in retrieval scenarios. Lightning IR provides a modular and extensible architecture that supports all stages of a retrieval pipeline: from fine-tuning and indexing to searching and re-ranking. Designed to be scalable and reproducible, Lightning IR is available as open-source: https://github.com/webis-de/lightning-ir.
@@ 1. Introduction
Pre-trained transformer-based language models have become a cornerstone in information retrieval (IR) research (Lin et al., 2022). Many different architectures have been proposed, each with their own implementation and training procedure. This plethora makes fine-tuning and comparing different model architectures cumbersome. However, all these models use similar backbones, are fine-tuned in the same way, and have only minor differences in their inference procedure.
To unify the usage of transformer-based language models in IR, we present the Lightning IR framework. Lightning IR builds on and extends PyTorch Lightning (Falcon and the PyTorch Lightning team, 2024) to provide several key features that set it apart from existing libraries for neural IR: (1) It is backbone agnostic, i.e., (almost) any HuggingFace (Wolf et al., 2020) transformer-based language model can be used. (2) It supports the entire IR pipeline, from fine-tuning and indexing to searching and re-ranking. (3) It is flexible and supports, for example, multi-vector or sparse bi-encoders and pointwise or listwise cross-encoders. (4) It provides an easy to use API and CLI. (5) It is highly configurable, allowing for reproducible experiments and painless model comparison. In this paper, we compare Lightning IR to existing frameworks, describe its features and API, and demonstrate Lightning IR’s capabilities.
@@ 2. Comparison to Similar Frameworks
Several existing frameworks support fine-tuning and inference with neural retrieval models, but they differ in the supported stages of the retrieval pipeline and the types of models they can handle (see Table 1 for an overview). Frameworks like RAGatouille, SentenceBERT (Reimers and Gurevych, 2019), or Seismic (Bruch et al., 2024) focus on specific model types, while frameworks like PyTerrier (Macdonald et al., 2021) or Experimaestro-IR (Piwowarski, 2020) support several model types, but not all are available for all stages.
Instead, Lightning IR is a framework that implements different models as generic, configurable, and extensible modules. In essence, a model is defined by its backbone encoder and how the contextualized embeddings of a query and document are combined to compute a relevance score. This design allows Lightning IR to support a wide range of model types for all stages. In addition, as all model types use the same API, they can use the same fine-tuning and inference procedures. This makes it easy to compare different model types and to experiment with new ones.
@@ 3.1. Model
Lightning IR supports two types of models: cross-encoders and bi-encoders. Both model types are built on top of backbone encoder models from HuggingFace (Wolf et al., 2020). For example, the following code shows how to initialize a new bi-encoder and cross-encoder model using some pre-trained model from HuggingFace, where {HF_MODEL} is a placeholder for the model name.
@@
A configuration class defines how each model type uses the contextualized embeddings generated by the backbone model to compute a relevance score. For example, a ColBERT-style model (Khattab and Zaharia, 2020) does not pool the contextualized embeddings. It instead uses late interaction to compute relevance scores over the contextualized query and document token embeddings. A ColBERT-style model can be initialized in Lightning IR as follows.
@@
For usability and reproducibility, we combine a Lightning IR model and tokenizer in a PyTorch Lightning module (Falcon and the PyTorch Lightning team, 2024). The module is responsible for handling training and inference logic, but it also provides convenience functions to quickly score queries and documents. For example, the following code snippet shows how to compute scores using a pre-trained BERT-based bi-encoder and an ELECTRA-based cross-encoder.
@@ 3.2. Dataset
Lightning IR tightly integrates with ir_datasets (MacAvaney et al., 2021) to provide access to a wide range of common information retrieval datasets, but custom datasets are also supported. Datasets are split into four different classes: document, query, tuple, and run datasets. Document datasets iterate over a document collection and are used for indexing. Query datasets iterate over queries and are used for retrieval. Tuple datasets are used for fine-tuning as they contain samples consisting of a query and multiple documents. Run datasets contain ranked documents for a query (optional: relevance judgments) and are used for re-ranking; also fine-tuning is possible by sampling n-tuples from rankings. The following code snippet shows how to load the MS MARCO passage dataset (Bajaj et al., 2018) and the TREC Deep Learning 2019 passage ranking dataset (Craswell et al., 2019).
@@
Multiple datasets can be combined into a single PyTorch Lightning datamodule. Similar to a model’s Lightning module, datamodules make fine-tuning and inference easier by handling data sampling and batching. They also cleanly separate fine-tuning and evaluation data. The following code snippet shows how to create a datamodule for fine-tuning a bi-encoder model on MS MARCO triples and evaluating it on the TREC Deep Learning 2019 and 2020 passage ranking datasets.
@@
RunDataset("msmarco-passage/trec-dl-2020"),
@@ 3.3. Trainer
To use an index for retrieval, the path of the index must be passed to a searcher class that matches the indexer class used to create the index. If a dataset has relevance judgments in ir_datasets and evaluation metrics are specified in the module’s configuration, the trainer will automatically evaluate the retrieval effectiveness. The following code snippet shows how to retrieve documents for a query using the indexed documents.
@@ 3.4. CLI
Lightning IR provides a command line interface (CLI) to simplify the usage of the framework. The CLI is built on top of PyTorch Lightning’s CLI and provides commands for fine-tuning, indexing, searching, and re-ranking. All options are configurable via command-line arguments or a configuration YAML file. The configuration YAML files are especially useful for reproducibility. For example, to fine-tune a bi-encoder model on the MS MARCO triples dataset, the following command can be used.
@@
The CLI also supports indexing, searching, and re-ranking via the index, search, and re_rank commands. The configuration file and the command for indexing the MS MARCO passage collection using a fine-tuned bi-encoder model are shown below. The configuration files for searching and re-ranking are similar.
@@ 4. Lightning IR: Supported Models
Lightning IR supports fine-tuning and running inference on a wide range of bi- and cross-encoder models but we have also added support for a number of popular models not natively fine-tuned in Lightning IR. This includes all bi-encoders from the sentence transformers library (Reimers and Gurevych, 2019), SPLADE models released by Naver labs (Formal et al., 2021), the official ColBERT checkpoints (Khattab and Zaharia, 2020), and monoT5 and RankT5 models (Nogueira et al., 2020; Zhuang et al., 2022). Further additional models can be easily added anytime by providing the corresponding configuration files and adding the model to the Lightning IR model registry.
@@
# class_path: BiEncoderModule # or CrossEncoderModule
@@
monoELECTRA Large (Schlatt et al., 2025) 0.750 0.791
@@ 5. Reproducibility Experiment
Table 3 compares the effectiveness of our fine-tuned models with the official checkpoints [^5] https://huggingface.co/sentence-transformers/msmarco-bert-base-dot-v5, [^6] https://huggingface.co/naver/splade-v3, [^7] https://huggingface.co/colbert-ir/colbertv2.0 provided by the authors on the TREC 2019 and 2020 Deep Learning track data (Craswell et al., 2019, 2020). Our fine-tuned models achieve competitive effectiveness compared to the official checkpoints. Minor differences between our results and the official checkpoints can be attributed to randomness in the training process and are not statistically significant, except for the SPLADE models (our SPLADE model is more effective on TREC Deep Learning 2019 and less effective on TREC Deep Learning 2020). These results demonstrate that reproducing the effectiveness of state-of-the-art models is possible with minimal effort in Lightning IR.
@@ 6. Conclusion
We have introduced Lightning IR, a PyTorch Lightning-based framework that enables straightforward fine-tuning and inference of transformer-based language models in retrieval tasks. By conducting a reproducibility experiment, we have demonstrated the capabilities of Lightning IR and the simplicity and flexibility of its API. With short code snippets and minimal effort, we were able to fine-tune and evaluate a variety of state-of-the-art models that achieve competitive effectiveness to their official checkpoints. Future work includes extending Lightning IR to support additional efficient dense, sparse, and multi-vector indexing and retrieval pipelines, such as PLAID (Santhanam et al., 2022) and Seismic (Bruch et al., 2024), and to improve the latency and scalability of multi-vector and sparse retrieval models.